# BoxingGym: Benchmarking Progress in Automated Experimental Design and Model Discovery

### v1

@@ Abstract
Understanding the world and explaining it with scientific theories is a central aspiration of artificial intelligence research. Proposing theories, designing experiments to test them, and then revising them based on data are fundamental to scientific discovery. Despite the significant promise of LLM-based scientific agents, no benchmarks systematically test LLM’s ability to propose scientific models, collect experimental data, and revise them in light of new data. We introduce BoxingGym , a benchmark with 10 environments for systematically evaluating both experimental design (e.g., collecting data to test a scientific theory) and model discovery (e.g., proposing and revising scientific theories). To enable tractable and quantitative evaluation, we implement each environment as a generative probabilistic model with which a scientific agent can run interactive experiments. These probabilistic models are drawn from various real-world scientific domains ranging from psychology to ecology. To quantitatively evaluate a scientific agent’s ability to collect informative experimental data, we compute the expected information gain (EIG), an information-theoretic quantity which measures how much an experiment reduces uncertainty about the parameters of a generative model. A good scientific theory is a concise and predictive explanation. Therefore, to quantitatively evaluate model discovery, we ask a scientific agent to explain their model and then assess whether this explanation enables another scientific agent to make reliable predictions about this environment. In addition to this explanation-based evaluation, we compute standard model evaluation metrics such as prediction errors. We find that current LLMs, such as GPT-4o, struggle with both experimental design and model discovery. We find that augmenting the LLM-based agent with an explicit statistical model does not reliably improve these results. [^1] Project: https://github.com/kanishkg/boxing-gym/tree/v0.1.0-beta
@@ 1 Introduction
Helping humans understand the world (and themselves) by discovering scientific theories has been a foundational goal of artificial intelligence research [25]. Proposing theories about the world, conducting experiments to test them, and revising them based on evidence is central to this process [6]. Recent advances in artificial intelligence, particularly with large language models (LLMs), have shown promising potential for accelerating scientific discovery. LLMs have extensive scientific knowledge [2], strong inductive reasoning capabilities [49, 38], and the ability to write statistical models of data [21, 22]. These promising results suggest that LLMs, functioning as autonomous agents, could be well-suited for two key components within the scientific discovery pipeline: experimental design (i.e., collecting informative experiments to test scientific theories) and model discovery (i.e., developing interpretable, parsimonious models based on experimental data).
Previous work has evaluated automated experimental design and model discovery in isolation [14, 15, 13, 21]. However, in realistic settings, they are tightly coupled: scientists collect experimental data to build better models and better models help scientists design more informative experiments. While autonomous scientific agents are an exciting prospect, there is currently no systematic way to evaluate an agent’s ability to propose scientific models, collect experimental data, and revise them in light of new data. This motivates the need for a benchmark that evaluates an agent’s capabilities within an integrated scientific discovery pipeline.
We now outline the key desiderata for a framework for evaluating experimental design and model discovery: (1) Our framework should enable the agent to actively experiment with the environment without requiring the agent to perform time-consuming and resource-intensive real-world lab experiments. (2) Since scientific theories come in different forms, our framework should flexibly accommodate different representations of scientific theories. (3) Our framework should evaluate experimental design and model discovery in an integrated way. (4) Science is often goal-directed or driven by a particular inquiry. For example, a biologist might perform experiments with the goal of identifying cellular mechanisms underlying circadian rhythm in mammals. Our framework should allow users to specify high-level goals to guide the agent’s discovery process. Our desiderata are inspired by the conceptual framework for scientific modeling introduced by George Box [8, 7], which emphasizes an iterative process of building models, designing experiments to test them, and revising models based on the results.
@@
We introduce principled evaluation metrics that measure the quality of experiments and discovered models. To evaluate experimental design, we draw from Bayesian optimal experimental (BOED) design [39] and use expected information gain (EIG) to measure the informativeness of an experiment. EIG captures how much an experiment reduces uncertainty in the parameters of a generative model and, importantly, this measure complements our decision to implement environments as generative models. To evaluate model discovery, we take inspiration from the fact that science is a communicative endeavor. We propose a communication-based evaluation strategy: we ask a scientist agent to distill their experiments into a natural language explanation and evaluate how much that explanation empowers a novice agent, who does not have access to the experiments conducted by the scientist, to make accurate predictions about the environment. Finally, as initial steps, we implement and evaluate two baseline agents: (1) a purely language-based agent (2) a language agent that builds statistical models of the data and uses the model to aid their predictions [21, 22]. These tasks are challenging for both agents and highlight fertile ground for further research.
@@ 2 Related Works
Bayesian optimal experimental design (BOED) is a principled framework for designing maximally informative experiments with successful applications in a wide variety of disciplines [45, 10, 31]. Although the framework is appealing, implementing BOED in practice is challenging because it involves computing information gain metrics, such as expected information gain (EIG), which are often intractable. There are a number of approaches to approximating the EIG of a design. Rainforth et al. [39] propose a nested Monte Carlo estimator and Foster et al. [14] estimate the EIG using a variational approximation. Subsequent work has scaled BOED to sequential settings by learning an amortized design network [15]. While these methods are significant technical innovations, they typically assume that the data is well-approximated by an assumed generative model. However, in many practical settings, we also want to revise our model as we collect new data.
@@
Automated model discovery from data has been a long-standing goal in AI, with the aim of building interpretable and explainable models that capture underlying patterns and relationships in data. Researchers have developed systems for discovering physical laws [5, 26], performing nonparametric regression [13], and conducting unsupervised learning [18]. Recently, Li et al. [21, 22] proposed integrating language models into an automated statistical model discovery pipeline that involves both proposing candidate models and critiquing them, leveraging their ability to generate symbolic expressions and critique their own reasoning. This work highlights the potential of using language models as a powerful tool for model discovery.
Reasoning with LLMs.
Experimental design and model discovery require both deductive reasoning (e.g., what does a given hypothesis entail) and inductive reasoning, (e.g., what might be the underlying hypothesis that explains the data) (c.f., Liu et al. [23]). Wang et al. [49], Qiu et al. [38] show how LLMs can reason inductively to solve abstract pattern based problems from a few examples. Other work in deductive reasoning has shown how LLMs [43, 42, 37] can perform logical deduction and identify inconsistencies or contradictions in a set of statements. However, these works have primarily been applied to deterministic, rule-based systems rather than the noisy, stochastic data typically seen in scientific discovery.
In-context Exploration.
The challenge of building accurate models through experimental design requires efficient exploration and information-seeking strategies. Recent research has tried to study this through in-context exploration — Nie et al. [32] and Min et al. [27] demonstrate how to optimize in-context exploration, while Monea et al. [29] shows how transformers can learn reinforcement learning policies directly in a sequence modelling framework. This work on exploration in language has been complemented by recent research on teaching language models how to search effectively in language [17, 16, 44, 20].
@@
BoxingGym takes inspiration from the long-line of research in designing RL environments [9, 28]. In particular, we adapt the modularity and simplicity of these benchmarks. However, the focus of our environments is on evaluation and not training agents. More recently, Jimenez et al. [19] presented an interactive benchmark to test capabilities of language-agents on debugging software issues. Our work builds on this line of research by introducing a benchmark for evaluating the ability of language agents to perform iterative model discovery through a combination of reasoning and experimentation.
@@ 3 Boxing Gym
We formalize the experimental design and model discovery problem within the framework of probabilistic modelling and Bayesian optimal experimental design (BOED). In BoxingGym , each environment is implemented as a generative model defining a joint distribution over the experimental outcome $y$, experimental design $d$, and unobserved parameters $\theta$. This joint distribution is characterized in terms of a prior distribution over $\theta$, $p(\theta)$ and a simulator $p(y|\theta,d)$ which is a model of the experimental outcome $y$ given parameters $\theta$ and design $d$. For example, $\theta$ could be the parameters of some behavioral model of participants in a psychology experiment, $d$ could be the questions posed to the participants, and $y$ could be the participant’s response to $d$. Within this framework, running an experiment corresponds to choosing a design $d$ and observing a sample $y$ from the marginal predictive distribution conditioned on that design, i.e., $y\sim p(y|d)=E_{p(\theta)}[p(y|\theta,d)]$)  [^2] In the sequential setting, we replace the prior $p(\theta)$ with the posterior $p(\theta|y,d)$..
@@ 3.2 Evaluation
We now discuss how we quantitatively evaluate a scientific agent’s capabilities in 1) experimental design and 2) model discovery.
@@ 3.2.1 Evaluating experimental design via Expected Information Gain
To quantitatively evaluate experimental design, we take inspiration from the Bayesian OED literature [14, 15]. Crucially, our design choice to implement environments as generative models enables us to leverage this literature. For each domain, we have an underlying predictive model $p(y|\theta,d)$. We quantify the informativeness of a design $d$ through the expected information gain (EIG), an information-theoretic quantity measuring the (expected) reduction in uncertainty about the model parameters $\theta$ after running an experiment $d$.
@@
[…]
@@ 3.2.2 Evaluating model discovery via communication
To evaluate the quality of a model, we employ both a number of standard model evaluation metrics (e.g., prediction MSE) as well as a metric loosely inspired by communication through a noisy channel, that takes advantage of the natural language interface. In particular, a scientist agent interacts with an environment via experiments. After a series of these experiments, we ask the scientist agent to synthesize their findings through an explanation. We then evaluate the extent to which that explanation empowers a novice agent to make accurate predictions about the environment without any interaction. Since a good explanation is both predictive and parsimonious, we set a token limit on the explanation. Crucially, this evaluation method can accommodate different forms of scientific theories. For example, in our experiments, we ask the scientist agent to produce a statistical model and then interpret and distill the model into a natural language explanation to guide the novice agent.
@@ 3.2.3 Evaluating goals via prediction
To evaluate success at achieving a specific goal (e.g., how do the populations of predator and prey change with time) we employ a prediction target (e.g., predict the population of predators at a particular time) and calculate a standardized prediction error. First, we compute the error between the predicted and true values. Then, we standardize this error with respect to the prior predictive mean, which is obtained by assuming a uniform prior over the design space. Specifically, for each domain, we sample a design $d$ uniformly from the design space and a parameter $\theta$ from the prior distribution $p(\theta)$. We then generate samples from the predictive model $p(y|\theta,d)$ and average over multiple $d$ and $\theta$ to obtain the prior predictive mean $\mu_{0}$ and variance $\sigma_{0}$. Let $\{y_{i}\}_{i=1}^{n}$ be the ground truth outputs for inputs $\{x_{i}\}_{i=1}^{n}$. and let $\{\hat{y_{i}}\}_{i=1}^{n}$ be the predictions of the agent. The standardized prediction error is then calculated using these quantities, providing a measure of the agent’s performance relative to the prior predictive mean. We use a domain-specific function $f$ computing the discrepancy between a prediction $\hat{y_{i}}$ and ground truth value $y_{i}$ (e.g., MSE). We compute the errors $\epsilon_{i}=f(\hat{y_{i}},y_{i})$ and $\epsilon_{\mu_{0}}=f(\mu_{0},y_{i})$. Finally, we compute the standardized error as $\frac{\epsilon_{i}-\epsilon_{\mu_{0}}}{\sigma{0}}$.
@@ 3.3 Design Decisions in Constructing BoxingGym
Our environments are grounded in real-world scientific models from various domains, ensuring the benchmark tests the agent’s ability to handle realistic scenarios. We implement these environment as pymc generative models to make active experimentation tractable.
@@
Each environment has a specific goal or an intent of inquiry, mirroring the nature of human scientific research. This encourages the agent to engage in targeted experimentation.
@@
We use a language-based interface for our experiments which enjoys several compelling advantages. First, a language-based interface is flexible and general, since scientific domains can be described using natural language. Second, language is a natural choice, given that our work is motivated by the promise of LLM-based systems. Third, a language-based interface is interpretable to humans.
@@
BoxingGym places a strong emphasis on measuring the quality of the agent’s discoveries through the explanations it provides (§ 3.2.2). This design decision is motivated by two considerations. From a theoretical perspective, science is fundamentally about developing better theories, and scientific theories are explanations of observed phenomena. From a practical perspective, communicating findings to the broader scientific community is an essential aspect of scientific research. By using language, we can avoid committing to a particular representation of a scientific theory. We illustrate this flexibility, by showing how different representations can be easily integrated within our method for measuring natural language explanations.
@@ 3.4 Domains
[15] The scientist observes a participant’s choices for different immediate rewards ($ir$), delayed rewards ($dr$), and delay periods ($D$ days) Fig. 2 (right). Goals include predicting choices of a participant or discount factors.
@@
[15] A disease spreads at a certain infection rate. The scientist can measure the number of infected individuals at different points of time to predict future infections or the infection rate.
@@
4 Experiments with Baseline Agents
Table 1: Performance of GPT-4o Across Different Tasks. Numbers shown are standardized errors. Errors are averaged across 5 runs.
Environment Goal Before After After
Experiments 10 Experiments Communication
Hyperbolic Discounting Choice 0.31 0.74 0.74
Discount -0.06 -0.06 -
Location Finding Signal 0.96 1.24 0.97
Source Location 1.29 -0.15 -
Death Process Num Infected 1.19 0.46 0.75
Infection Rate 0.13 1.64 -
IRT Correctness 0.00 0.00 -0.28
Dugongs Length 0.06 -0.09 -0.08
Peregrines Population 2.29 -0.65 -0.63
Mastectomy Survival 0.18 0.27 1.00
Predator-Prey Population 0.08 -0.45 -0.26
Emotions Prediction 0.74 0.82 0.87
Moral Machines Judgement 0.32 0.44 0.60
Table 2: Performance of Box’s Apprentice Across Different Tasks. Standardized errors shown here. Errors are averaged across 5 runs.
Environment Goal Before After After
Experiments 10 Experiments Communication
Hyperbolic Discounting Choice 0.66 1.17 0.66
Location Finding Signal 0.99 1.45 1.18
Death Process Num Infected 3.79 -1.02 0.58
IRT Correctness 0.44 -0.12 -0.08
Dugongs Length 0.26 -0.08 -0.09
Peregrines Population 2.71 0.04 0.97
Mastectomy Survival 0.14 0.55 0.91
Moral Machines Judgement 0.97 0.89 0.56

### v2

# Scaling Environments for Agents (SEA)

@@ Abstract
Understanding the world and explaining it with scientific theories is a central aspiration of artificial intelligence research. Proposing theories, designing experiments to test them, and then revising them based on data are key to scientific discovery. Despite the promise of LLM-based scientific agents, no benchmarks systematically test their ability to propose scientific models, collect experimental data, and revise them in light of new data. We introduce BoxingGym , a benchmark with 10 environments for evaluating experimental design (e.g., collecting data to test a scientific theory) and model discovery (e.g., proposing and revising scientific theories). To enable quantitative and principled evaluation, we implement each environment as a generative probabilistic model with which a scientific agent can run interactive experiments. These probabilistic models are drawn from various real-world scientific domains ranging from psychology to ecology. To evaluate a scientific agent’s ability to collect informative experimental data, we compute the expected information gain (EIG), an information-theoretic quantity which measures how much an experiment reduces uncertainty about the parameters of a generative model. A good scientific theory is a concise and predictive explanation. To quantitatively evaluate model discovery, we ask a scientific agent to explain their model and evaluate whether this explanation helps another scientific agent make more accurate predictions. We evaluate several open and closed-source language models of varying sizes. We find that larger models (32B) consistently outperform smaller variants (7B), and that closed-source models generally achieve better results than open-source alternatives. However, all current approaches struggle with both experimental design and model discovery, highlighting these as promising directions for future research.  [^1] Project: https://github.com/kanishkg/boxing-gym
@@ 1 Introduction
Helping humans understand the world (and themselves) by discovering scientific theories is a foundational goal of artificial intelligence research [31]. Proposing theories about the world, conducting experiments to test them, and revising them based on data is central to this process [9]. Recent advances in large language models (LLMs), have shown promising potential for accelerating scientific discovery. LLMs have extensive scientific knowledge [2], strong inductive reasoning capabilities [53, 43], and the ability to propose models of data [27, 28, 11]. These promising results suggest that LLMs, functioning as autonomous agents, could be well-suited for experimental design (i.e., collecting informative experiments to test scientific theories) and model discovery (i.e., developing interpretable models based on experimental data).
Previous work has evaluated automated experimental design and model discovery in isolation [16, 17, 15, 27]. However, they are fundamentally coupled in real-world settings: scientists collect experimental data to build better models and better models inform better experiments. While scientific agents are promising, there is currently no systematic way to evaluate an agent’s ability to propose scientific models, collect experimental data, and revise them in light of new data. This motivates the need for a benchmark that evaluates an agent’s capabilities holistically in an integrated scientific discovery pipeline.
We outline the key desiderata for a framework that evaluates experimental design and model discovery: (1) The framework should enable the agent to actively experiment with the environment without requiring the agent to perform time-consuming and resource-intensive real-world lab experiments. (2) Since scientific theories come in different forms, the framework should flexibly accommodate different representations of scientific theories. (3) The framework should evaluate experimental design and model discovery in an integrated way. (4) Science is often goal-directed or driven by an inquiry. For example, a biologist might perform experiments with the goal of identifying cellular mechanisms underlying circadian rhythm in mammals. Our framework should allow users to specify high-level goals to guide the agent’s discovery process. Our desiderata are inspired by the framework for scientific modeling introduced by George Box [7, 8], which emphasizes an iterative process of building models, designing experiments to test them, and revising them accordingly.
@@
We introduce principled evaluation metrics that measure the quality of experiments and discovered models. To evaluate experimental design, we draw from Bayesian optimal experimental (BOED) design [44] and use expected information gain (EIG) to measure the informativeness of an experiment. EIG captures how much an experiment reduces uncertainty in the parameters of a generative model and, importantly, this measure complements our decision to implement environments as generative models. To evaluate model discovery, we take inspiration from the fact that science is a communicative endeavor. We propose a communication-based evaluation strategy: we ask a scientist agent to distill their experiments into a natural language explanation and evaluate how much that explanation empowers a novice agent, who does not have access to the experiments conducted by the scientist, to make accurate predictions about the environment.
We evaluate several open and closed-source language models ranging from 7B to 32B parameters. We find that larger models consistently outperform smaller variants, and closed-source models generally achieve better results than open-source alternatives. We also evaluate Box’s Apprentice [27], which augments language models with statistical modeling capabilities, but find that this augmentation does not reliably improve performance. Notably, we observe substantial variation in difficulty across environments, which remaining challenging even for the strongest models. Promisingly, some environments show clear performance improvements with model scale. These results highlight significant opportunities for improving automated scientific reasoning.
@@ 2 Related Works
Bayesian optimal experimental design (BOED) is a principled framework for designing maximally informative experiments across various disciplines [49, 12, 35]. While theoretically appealing, BOED’s practical implementation is challenging due to the intractability of information gain metrics like expected information gain (EIG). Although several methods [44, 16, 17] exist to approximate EIG, they assume the data follows a fixed generative model—limiting their utility when model revision is needed as new data is collected.
@@
Automated model discovery from data has been a long-standing goal in AI, aiming to build interpretable models that capture underlying patterns in data—from physical laws [6, 32] to nonparametric regression [15]. Recent work [27, 28] has integrated language models into this process, leveraging their ability to both propose and critique candidate models, demonstrating their potential as tools for automated model discovery. This work highlights the potential of using language models as a powerful tool for model discovery.
Reasoning and Exploration with LLMs.
Language models have shown promising capabilities in both deductive reasoning (deriving consequences from hypotheses) [47, 46, 42] and inductive reasoning (inferring hypotheses from observations) [53, 43]. While reinforcement learning has improved LLMs’ reasoning abilities [24, 21, 20, 22], these advances have primarily focused on deterministic, verfiable systems rather than the stochastic data typical in scientific discovery. Efficient exploration and information-seeking are crucial for experimental design and model building. Recent work [37, 33, 19, 18, 48, 26] has investigated in-context exploration strategies and shown how language models can learn how to search and explore directly through sequence modeling, developing effective search strategies in language.
@@
Drawing inspiration from established reinforcement learning principles [10, 34], BoxingGym adopts the modularity and simplicity of classic environments like OpenAI Gym while shifting focus to evaluation rather than agent training. While recent work has expanded interactive benchmarks to language agents —spanning tasks from software debugging [25] to automated scientific research[36, 29], our work advances this direction by introducing a principled framework for evaluating language agents’ capabilities in iterative experimental design and model discovery.
@@ 3 Boxing Gym
We formalize experimental design and model discovery using probabilistic modeling and Bayesian optimal experimental design (BOED). In BoxingGym , each environment is implemented as a generative model defining a joint distribution over the experimental outcome $y$, experimental design $d$, and unobserved parameters $\theta$. This joint distribution is defined in terms of a prior distribution over $\theta$, $p(\theta)$ and a simulator $p(y|\theta,d)$ which is a model of the experimental outcome $y$ given parameters $\theta$ and design $d$. For example, in a psychology experiment, $\theta$ could be the parameters of a behavioral model of participants, $d$ could be the questions posed to participants, and $y$ could be the participant’s response to $d$. Running an experiment corresponds to choosing a design $d$ and observing a sample $y$ from the marginal predictive distribution conditioned on that design, i.e., $y\sim p(y|d)=E_{p(\theta)}[p(y|\theta,d)]$)  [^2] In the sequential setting, we replace the prior $p(\theta)$ with the posterior $p(\theta|y,d)$..
@@ 3.2.1 Evaluating experimental design via Expected Information Gain
To evaluate experimental design, we take inspiration from the Bayesian OED literature [16, 17]. Crucially, our choice to implement environments as generative models enables us to leverage this literature. For each domain, we have an underlying predictive model $p(y|\theta,d)$. We quantify the informativeness of a design $d$ through the expected information gain (EIG), that measures the reduction in posterior uncertainty about the model parameters $\theta$ after running an experiment $d$. Below, $H$ is the Shannon entropy.
@@
Since the EIG is typically not available in closed-form, we use a Nested Monte Carlo estimator $\hat{\mu}_{\text{NMC}}(d)=\frac{1}{N}\sum_{n=1}^{N}\log\left(\frac{p(y_{n}|\theta_{n,0},d)}{\frac{1}{M}\sum_{m=1}^{M}p(y_{n}|\theta_{n,m},d)}\right)\quad\text{where}\quad\theta_{n,m}\overset{\text{i.i.d.}}{\sim}p(\theta),\;y_{n}\sim p(y|\theta=\theta_{n,0},d)$ We chose this estimator because it is a consistent estimator of the true EIG [44] and is straightforward to implement. EIG measures the value of an experiment under the assumption that the true distribution of experimental outcomes is modeled by $p(y|d)$. In general, this assumption is not true, but EIG is still a useful measure since we generate data from an underlying model in our benchmarks.
@@ 3.2.2 Evaluating model discovery via communication
To evaluate the quality of a model, we use standard model evaluation metrics (e.g., prediction MSE) and a communication-based metric that takes advantage of the natural language interface. In particular, a scientist agent interacts with an environment through experiments. After these experiments, we ask the scientist agent to synthesize their findings through an explanation. We then evaluate how much that explanation enables a novice agent to make more accurate predictions about the environment without any additional experiments. Since a good explanation is both predictive and parsimonious, we set a token limit on the explanation. Crucially, this evaluation method can accommodate different forms of scientific theories. In our experiments, we ask the scientist agent to produce a statistical model and then distill the model into a natural language explanation to guide the novice agent.
@@ 3.2.3 Evaluating goals via prediction
To evaluate success at achieving a specific goal (e.g., how do the populations of predator and prey change with time) we employ a prediction target (e.g., predict the population of predators at a particular time) and calculate a standardized prediction error. First, we compute the error between the predicted and true values. Then, we standardize this error with respect to the prior predictive mean, which is obtained by assuming a uniform prior over the design space. Specifically, for each domain, we sample a design $d$ uniformly from the design space and a parameter $\theta$ from the prior distribution $p(\theta)$. We then generate samples from the predictive model $p(y|\theta,d)$ and average over multiple $d$ and $\theta$ to obtain the prior predictive mean $\mu_{0}$ and variance $\sigma_{0}$. Let $\{y_{i}\}_{i=1}^{n}$ be the ground truth outputs for inputs $\{x_{i}\}_{i=1}^{n}$. and let $\{\hat{y_{i}}\}_{i=1}^{n}$ be the predictions of the agent. The standardized prediction error is then calculated using these quantities, providing a measure of the agent’s performance relative to the prior predictive mean. We use a domain-specific function $f$ computing the discrepancy between a prediction $\hat{y_{i}}$ and ground truth value $y_{i}$ (e.g., MSE). We compute the errors $\epsilon_{i}=f(\hat{y_{i}},y_{i})$ and $\epsilon_{\mu_{0}}=f(\mu_{0},y_{i})$. Finally, we compute the standardized error as $\frac{\epsilon_{i}-\epsilon_{\mu_{0}}}{\sigma_{0}}$. Crucially, since this metric is computed with respect to the prior predictive, this metric can be negative.
@@ 3.3 Design Decisions in Constructing BoxingGym
Our environments are grounded in real-world scientific models from several domains, ensuring the benchmark tests the agent’s ability to handle realistic scenarios. We implement these environment as pymc generative models to make active experimentation an automatic and tractable process.
@@
Each environment has a specific goal, mirroring the inquiry-driven nature of scientific research. This encourages the agent to engage in targeted experimentation.
@@
We use a language-based interface for our experiments because it’s flexible (i.e., scientific domains can generally be described in language), easily integrates with LLMs, and interpretable to humans.
@@
BoxingGym places a strong emphasis on measuring the quality of the agent’s discoveries through the explanations it can provide after experimentation (§ 3.2.2). This design decision is motivated by two considerations. From a theoretical perspective, science is fundamentally about developing better theories, and scientific theories are explanations of observed phenomena. From a practical perspective, communicating findings to the broader scientific community is an essential aspect of scientific research. By using language, we do not have to commit to a particular representation of a scientific theory. We illustrate this flexibility, by showing how different representations can be easily integrated within our method for measuring natural language explanations.
@@ 3.4 Domains
[17] The scientist observes a participant’s choices for different immediate rewards (irir), delayed rewards (drdr), and delay periods ($D$ days) Fig. 2 (right). Goals include predicting choices of a participant or discount factors.
@@
[17] A disease spreads at an infection rate. The scientist can measure the number of infected individuals at different points of time to predict future infections or the infection rate.
@@
4 Experiments